### v2

# Towards Efficient and Accurate Detection of On-Chip Fail-Slow Failures for Many-Core Accelerators

Abstract
Many-core accelerators are essential for high-performance deep learning, but their performance is undermined by widespread fail-slow failures. Detecting such failures on-chip is challenging, as prior methods from distributed systems are unsuitable due to strict memory limits and their inability to track failures across the hardware topology.
[…]
I Introduction
Many-core accelerators, integrating tens to thousands of processing cores on a single chip or via multi-chiplet packaging, have emerged as the backbone of modern applications. To support massive parallelism, these architectures rely on a tightly coupled design, where cores are interconnected via a high-bandwidth Network-on-Chip (NoC) to enable frequent, low-latency communication. This fabric allows them to efficiently handle throughput-demanding workloads [83, 14, 78, 13, 80], ranging from large-scale model training and inference to graph analytics [17, 64, 11] and scientific simulations [50, 69, 2]. As a result, many-core accelerators are increasingly deployed across diverse platforms from edge devices to data centers, delivering significant gains in performance and energy efficiency compared to traditional multi-core CPUs and GPGPUs [44, 35, 73, 74].
(a) Fail-slow failures on different hardware components.
(b) Normalized execution time of ResNet-50.
[…]
The characteristics of fail-slow failures, especially the transiency, mandate an on-chip detection paradigm, as traditional off-chip mechanisms are fundamentally ill-suited to capture such phenomena. On the one hand, the transient nature of these failures creates a critical detection gap for external observers. Off-chip monitoring typically relies on periodic polling or remote telemetry, operating at a coarse granularity and with significant transmission latency. By the time the status data reaches an external host, the micro-architectural stall effectively vanishes, rendering the external detector blind to the fleeting root cause. On the other hand, precise diagnosis requires high-resolution traffic data, which inevitably collides with the I/O bandwidth wall. Streaming continuous, fine-grained execution traces off-chip imposes a prohibitive bandwidth overhead. This heavy data movement not only effectively saturates limited I/O resources but also induces a parasitic effect: the diagnosis traffic competes with the actual workload for memory access, distorting the very system behavior it attempts to measure. Therefore, effective identification necessitates a shift to on-chip detection. Crucially, to accommodate hardware constraints, this approach must be lightweight and capable of performing in-situ statistical aggregation, capturing the failure root cause locally without the latency or bandwidth penalties of external analysis.
TABLE I: Comparison of hardware failure detection approaches in different scenes.
Target Scenario Detection Approach Failure Pattern Failure Model Validation Method Memory Overhead
Cloud System Sieve[15] Fail-Slow Latency Simulation 8 GiB
Micro Service Groot[71] Fail-Stop Fail-Slow Error Degradation Real Plat. 20 GiB
Sleuth[21] Fail-Slow Contention Simulation Real Plat. 84 MiB
Sage[20] Fail-Slow Latency Contention Real Plat. Not Mentioned
Distributed System IASO[55] Fail-Slow Timeout Real Plat. 1.6 MiB Per Hour
Perseus[46] Fail-Slow Latency Real Plat. Not Mentioned
ADR[45] Fail-Slow Packet Loss Latency Real Plat. Not Mentioned
Many-Core Accelerator SLOTH (Ours) Fail-Slow Degradation Simulation 150 KiB
While fail-slow detection has been explored in microservices and distributed systems (Table I), adapting these paradigms to many-core accelerators is impeded by two challenges. First, many-core systems have strict on-chip memory limits. Standard runtime tracing strategies [55, 46] and graph-based analyses [71, 21, 33] rely on capturing extensive execution logs, incurring massive storage overheads (e.g., Gigabytes for Sieve and Groot) that contradict the hardware reality shown in Figure 2a. Even a minimal 1-second trace of DarkNet-19 necessitates $2\text{\,}\mathrm{MiB}$ just for communication logs; this creates an unbridgeable gap with on-chip SRAM capacity (typically $64\text{\,}\mathrm{KiB}$ per core), rendering continuous in-situ tracing spatially infeasible. Second, existing methods are lack of hardware topological awareness. while graph-based methods analyze software dependencies, they struggle to distinguish root causes from propagated symptoms due to the complex on-chip interconnect topology. As illustrated in Figure 2b, a fail-slow failure in Core 1 triggers a cascading slowdown: the delayed data transmission propagates along the communication links, forcing downstream cores (Core 2 and Core 3) into involuntary wait states. Because existing methods focus on application-level stragglers without accounting for the underlying hardware backpressure dynamics, they inevitably conflate the symptoms with the cause, misidentifying the backpressured victims as the culprits.
Figure 2: Challenges for detecting fail-slow failures: (a) On-chip, off-chip memory capacity versus computation and communication tracing log size for tracing DarkNet-19 for 1 second on a $\times4$ core processor. (b) Slowdown propagates across cores along communication dependencies.
Therefore, an effective detection mechanism for many-core accelerator must be both lightweight to be efficient and hardware-aware to be accurate. A lightweight design must operate within tight memory limits, while a hardware-aware approach must analyze the accelerator’s hardware topology to localize the root cause.
In this paper, we present SLOTH [^1] The source code of SLOTH along with failure datasets and traces are available at https://anonymous.4open.science/r/sloth-91D7(Slowdown Localization via On-chip Tracing and Hierarchical graph analysis), a fail-slow detection framework with a comprehensive workflow that spans data collection, compression, storage, and analysis to meet these requirements.
First, we propose a workload-aware probe mechanism that enables lightweight performance data tracing without interfering with program execution. It automatically analyzes the computation graph to identify insertion points, collecting operator-level performance metrics. We further design a sketch-based data structure to compress and store the collected traces efficiently. To narrow down the failure candidates, we employ a group-based fail-core detection in combination with underdetermined linear system-based fail-link inference. To capture both software and hardware propagation of fail-slow failures, we reconstruct the communication dependency graph onto the hardware topology, and incorporate temporal information through hierarchical extension. Finally, we introduce FailRank, an iterative ranking algorithm that computes a fail-slow score for each component, thereby identifying the most likely root cause of fail-slow failures. In summary, our contributions are:
  • SL-Compiler: A workload-aware probe mechanism that leverages compiler-based instrumentation to automatically insert probe at operator boundaries. This enables operator-level tracing with minimal overhead while capturing fine-grained metrics.
  • SL-Recorder: A compact, on-the-fly, sketch-based data structure that filters and compresses trace data. This design preserves critical failure signatures while operating within the strict memory constraints of on-chip SRAM.
  • SL-Tracer: A topology-aware, multi-stage diagnosis pipeline that integrates the software communication graph with the hardware topology. It uses a novel algorithm—FailRank—to accurately trace failure propagation across cores and links, and pinpoint the true root cause.
[…]
II Background
II-A Many-Core DNN Accelerators
Many-core DNN accelerators integrate numerous cores into a scalable on-chip network to exploit massive parallelism for deep learning workloads. Figure 3 illustrates the fundamental components of such an architecture. Inter-core communication is enabled by the Network-on-Chip (NoC), which can adopt various topologies. Common structures include 2D-mesh, torus, and ring networks [83, 18], while more complex topologies such as hypercube and dragonfly have also been proposed [19]. These topologies can support different communication patterns and, in specific scenarios, provide higher bandwidth. In this work, we focus on mesh-based networks, as they are implementation-friendly and are the most widely adopted topology in practice [68, 22]. Each core is directly connected to a router and consists of three major components: a control unit, a compute unit, and a scratchpad memory. The control unit manages intra-core resource allocation and task scheduling. The compute unit executes DNN operators and typically integrates specialized accelerators such as vector units or tensor units optimized for matrix operations. The scratchpad memory is a small, high-speed SRAM placed adjacent to the compute unit, typically less than $64\text{\,}\mathrm{KiB}$ in capacity, storing activations, weights, and intermediate results. It is explicitly managed by software or compilers, thereby simplifying hardware coherence logic [16, 4].
Figure 3: Overview of a many-core DNN accelerator with 2D-mesh network and its intra-tile organization.
Based on the hardware organization above, many-core DNN accelerators employ a dataflow execution model to maximize concurrency and overlap computation with communication[9, 28, 7]. In this paradigm, DNN models are represented as computation graphs, where nodes denote operators (e.g., convolution, pooling) and edges represent data dependencies between layers[12, 54]. Execution is data-driven—an operator is triggered as soon as its input features and weights become available, avoiding the cost of global synchronization. DNN operators are mapped onto cores according to data locality and the layer-wise computational workflow, while the scheduler within each core monitors operator readiness, fetches operands from local or neighboring cores, and dispatches them to compute units. Intermediate features are routed through the NoC according to the communication pattern derived from the DNN computation graph. This fine-grained pipelining of data movement and operator processing improves hardware utilization, but also creates tight coupling between cores: delays in one core’s execution can propagate through dependent operators across the chip, amplifying the performance impact of fail-slow failures in DNN inference.
II-B Hardware Failures
In many-core DNN accelerators, hardware components are subject to a wide range of failures that invariably degrade inference performance. Based on their behavioral characteristics, common hardware failures are classified into two categories: fail-slow and fail-stop[23, 55].
For cores, bit flips can alter register values, leading to incorrect computation results or even program crashes [60, 27], which are characteristic of fail-stop failures. In contrast, performance variations caused by internal resource contention, workload imbalance, or hardware aging typically degrade overall performance [67, 85], and are categorized as fail-slow failures. For memory, permanent physical defects as well as software bugs or mismanagement can trigger system crashes, thus are fail-stop failures. Meanwhile, memory leaks or buffer overflows could consume significant system resources, causing response delays and performance degradation, which belong to fail-slow failures[47, 79]. For NoC, transient bandwidth reductions on links or low router port utilization can decrease overall throughput and prolong program execution time[8, 70], both of which are fail-slow failures. In contrast, routing-induced deadlocks can completely stall communication, making them fail-stop failures[66].
In this paper, we focus on detecting fail-slow failures caused by core-level performance variance and link-level bandwidth degradation—both of which are prevalent and, more importantly, can propagate along communication and execution dependencies, leading to system slowdowns.
Figure 4: Overview of SLOTH.
III The SLOTH Framework
SLOTH is an automated framework for detecting and locating fail-slow failures in many-core DNN accelerators. As shown in Figure 4, the framework is composed of three main components. SL-Compiler instruments DNN operators with lightweight probes at operator boundaries, collecting fine-grained runtime execution traces with minimal overhead. SL-Recorder then filters irrelevant traces and compresses redundant ones, producing a compact yet representative list of communication and execution patterns. Finally, SL-Tracer leverages the compressed data to construct multi-level communication graphs and performs the FailRank algorithm on them, pinpointing fail-slow cores or links responsible for performance degradation. Together, these components form a systematic pipeline that bridges low-overhead monitoring with high-accuracy failure diagnosis.
III-A Input and Output
As shown in the left of Figure 4, the inputs to SLOTH consist of: (1) DNN Workload: DNN models are composed of operators such as convolution, pooling, and data movement instructions including read, send, and recv. The framework supports arbitrary operator-to-core mappings [84, 36, 25]. In our implementation, we use a Gemini-like [7] mapping strategy for evaluation; (2) Arch Config: the specification of the target many-core DNN accelerator architecture, which enables analysis across different hardware platforms; (3) Probe Config: parameters that determine what execution information to collect, where to place probes at operator boundaries, and how to store the collected trace data; (4) Failure Model: definitions of fail-slow behavior, including failure duration, performance degradation rate, and affected components.
The outputs fall into two categories: fail-slow detection results, including the root causes, SL-Recorder’s optimal compression parameters, and the overhead of data collection and storage; and system-level metrics, such as total inference time and per-layer execution latency.
III-B SL-Compiler
Traditional tracing approaches face a trade-off between coverage and overhead: exhaustive collection is prohibitively expensive, while reactive collection risks missing critical failure signatures. To address this challenge, we propose SL-Compiler, a DNN-aware probe instrumentation tool that analyzes the hierarchical structure and explicit data dependencies of DNN workloads to insert eBPF (extended Berkeley Packet Filter) inspired probes, reducing monitoring overhead while preserving the ability to capture fail-slow behaviors.
TABLE II: Components of probe’s five-tuple definition.
Component Option Description
Fragment Exec Record calculation related traces
Route Record communication traces
Mem Record memory access traces
Type Comp Match compute instructions
Comm Match communication instructions
IO Match input / output instructions
Location Pre Before the target instruction
Post After the target instruction
Surround Before & after the instruction
Level Inst Aggregate trace at instruction level
Stage Aggregate trace at stage level
Structure List Store traces in list manner
Sketch Store traces using SL-Recorder
Probe Abstraction. SL-Compiler decomposes the monitoring process into three configurable abstractions, as summarized in Table 1. Probe fragments define what data to collect—whether computation time (Exec), communication latency (Route), or memory access (Mem). Insertion tracepoints specify where to place probes by combining an instruction type filter (Type: Comp, Comm, or IO) with a relative position (Location: Pre, Post, or Surround). Data maps determine how to organize collected traces, supporting both fine-grained (Level: Inst) and coarse-grained (Level: Stage) aggregation, with flexible storage structures (Structure: List for dense access or Sketch for memory efficiency). Together, these abstractions enable users to express complex monitoring strategies through a compact five-tuple configuration.
Figure 5: Examples of SL-Compiler probing.
DNN-graph-guided Insertion. A key innovation of SL-Compiler is its ability to automatically generate probe configurations by analyzing the DNN computation graph and its mapping to the target architecture. As illustrated in Figure 5, this process takes as input the workload mapping—which specifies how DNN operators are distributed across cores—and performs static analysis to identify critical monitoring points. The process proceeds in three steps. First, it parses the computation graph to extract the layer sequence, data dependencies, and operator types (e.g., Conv2D, Pooling). Second, it analyzes the workload-to-core mapping to determine which cores execute communication-intensive operations (requiring Route probes) versus computation-heavy operators (requiring Exec probes). For instance, in Figure 5b, it identifies that Conv layers dominate computation and automatically inserts post-execution probes ([Exec, Comp, Post, Stage, List]) to capture their execution time. Similarly, in Figure 5c, it detects inter-core data transfers and inserts pre-communication probes to monitor link latency. Third, it generates the instrumented code by inserting lightweight probe fragments at the identified tracepoints. Importantly, this process is fully automated, freeing users from manual instrumentation while still supporting custom probe specifications for specialized analysis needs.
Operator-level Semantic Collection. Beyond basic performance metrics, SL-Compiler captures high-level DNN operator semantics to facilitate root cause analysis. Each probe records not only the standard execution attributes—such as start/end timestamps, and core ID, but also DNN-specific metadata that characterizes the computational context. For computation probes, this includes the operator type (e.g., Conv2D, GEMM), input/output tensor dimensions, and floating-point operation count (FLOPs). For communication probes, it records the data volume transferred, the source and destination cores, and the corresponding tensor being communicated. This semantic information serves two critical purposes: (1) it enables layer-aware analysis in SL-Tracer, allowing the framework to correlate slowdowns with specific DNN layers; and (2) it facilitates communication pattern recognition, helping distinguish between normal inter-layer data transfers and anomalous traffic caused by fail-slow failures. For example, if a Conv layer exhibits unexpectedly high communication latency, the recorded tensor dimensions can reveal whether the slowdown stems from excessive feature map sizes or from degraded link bandwidth.
III-C SL-Recorder
To reduce the storage overhead, we propose SL-Recorder, a lightweight on-the-fly data structure for trace compression and pattern-based filtering. The data structure is also called Fail-Slow Sketch as it is inspired by sketches. As illustrated in Figure 6, Fail-Slow Sketch operates in two stages. In Stage-1, a Running Track identifies frequently occurring traces with similar execution behaviors. In Stage-2, the common patterns of the traces are extracted to reduce storage redundancy.
Input: instruction trace $e$; number of hash tables $d$; number of buckets $m$; pattern threshold $H$
1 for each $i\in[1,d]$ do
2  $Index\leftarrow\text{Hash}(e.key)\bmod m$;
3  $bucket\leftarrow\text{HashTable}[i][Index]$;
4
5 if bucket.key equals to e.key then
6     increase the frequency of e by 1;
7     if frequency of $e\geq H$ then Insert_Stage_2(e, attributes) ;
8
9 else if bucket is empty then
10     $bucket.key\leftarrow e.key$;
11     set the frequency of e to 1;
12
13 else
14     decrease the frequency of e by 1;
15     if frequency of e is 0 then clear bucketbucket to empty ;
16
17
18
19 Function Insert_Stage_2($e$, attr)
20  if $e.key\notin\text{Stage-2}$ then
21     if Stage-2.length $\geq$ MAX_LENGTH then evict the earliest pattern in Stage-2 ;
22     insert $e$ into Stage-2;
23     return 1;
24
25 else
26     Update($e.key$, attr);
27
28 return 0;
29
Algorithm 1 Fail-Slow Sketch Insertion
Stage-1 consists of $d$ bucket arrays, each associated with a distinct hash function. These hash functions map trace features that characterize execution patterns to bucket indices. Each array contains $m$ buckets, and each bucket has two fields: trace pattern (key) and its frequency. When the frequency exceeds a threshold $H$, the corresponding trace pattern is considered as a potential candidate for fail-slow pattern.
Stage-2 maintains a list of candidate fail-slow patterns and their associated attributes. This list supports operations such as updates and evictions, and its maximum length (MAX_LENGTH) is determined based on the characteristics of the workload. In this stage, traces with the same pattern are compressed, and statistical summaries are computed to support root cause analysis. When the list reaches its maximum capacity, the Fail-Slow Sketch evict the most “healthy” pattern to make room for newly observed patterns.
Insertion. Given an input trace $t$, Fail-Slow Sketch first extracts its pattern features and computes the target bucket index using corresponding hash function. The trace is then inserted into the bucket (donated as $b$). As shown in Algorithm 1, there are three possible cases:
Case 1: $t$ is not in the bucket, and $b$ is currently empty. In this case, the trace pattern is stored in $b$, the frequency is set to 1, and the insertion completes.
Case 2: $t$ is not in the bucket, and $b$ is occupied. In this case, the frequency of $b$ is decremented by 1. If the frequency drops to zero, the bucket is cleared, and the algorithm terminates.
Case 3: $t$ is in the bucket $b$. In this case, the frequency counter is incremented by 1. If the frequency reaches the threshold $H$, the trace pattern is considered a candidate for Stage-2. If the pattern already exists in the Stage-2 list, its statistics are updated via update function; otherwise, a new pattern entry will be created. If the Stage-2 list has reached its capacity, typical replacement strategies such as frequency-based, probability-based, or arrival-time-based policies can be applied. In our design, we use an arrival-time-based replacement policy to ensure correct insertion of trace patterns.
Figure 6: An example of Fail-Slow Sketch (d=1).
Update. Given a trace pattern and its attributes, the update operation modifies the corresponding entry in the Stage-2 pattern list to maintain the overall statistical information. For communication patterns, this includes updating metrics like data volume. For computation patterns, the update includes floating-point operation count, core id, and other relevant metrics. Additionally, all trace patterns have to maintain the timestamp of their first appearance and their duration, which are essential for root cause analysis.
Example: Figure 6 shows a running example of Fail-Slow Sketch. To insert coming trace $Comp4$, Stage-1 extracts its pattern $cp4$. $cp4$ has been inserted into corresponding bucket before, so Fail-Slow Sketch adds the frequency of the bucket by 1 and finishes inserting. When inserting trace $Comm2$, the frequency of the bucket reaches threshold $H=10$ and there is space in Stage-2’s pattern list. Therefore, Fail-Slow Sketch create a new trace pattern (cm2cm2,10,128,5) for $Comm2$ through Creation operation, in which $cm2$ is the trace pattern, 10 is $cm2$’s frequency, 128 is the data volume and 5 is the destination of communication. Besides, Fail-Slow Sketch performs Update operation for existing pattern $cm1$ and Eviction operation for deleted pattern $cp3$. After inserting all the traces, the compressed traces, also the trace patterns are stored in Stage-2’s pattern list.
Mathematical Analysis. It can be proved that increasing the number of hash functions $d$ can effectively improve the retention probability, but also increases computational overhead; increasing the number of buckets $m$ helps reduce hash collisions, but also increases storage overhead. We provide a proof for the conclusion as follow.
Given a fixed workload, let $N$ be the number of instruction traces, $f_{i}$ be the frequency of trace $t_{i}$, $B_{j}[k]$ be $k$-th bucket of $j$-th bucket array, and $F_{i,j,k}$ be the number of traces mapped to bucket $B_{j}[k]$ except for trace $t_{i}$.
LEMMA 3.1. For any trace pattern $R_{i}$ in the instruction traces, let $P(R_{i})$ be its retention probability in Stage-2, then
$P(R_{i})\geq 1-(\frac{N-f_{i}}{m(f_{i}-H)})^{d}$
PROOF. For a single bucket $A_{j}[k]$ in Stage-1 which records trace $t_{i}$, according to the insertion algorithm, we have
$f_{i}-F_{i,j,k}\leq A_{j}[k].freq\leq f_{i}$
Since the hash function is uniformly distributed, the probability of mapping to a specific bucket is $\frac{1}{m}$, so the expectation of $F_{i,j,k}$ is
$E[F_{i,j,k}]=\Sigma_{t\neq t_{i}}f_{t}\frac{1}{m}=\frac{N-f_{i}}{m}$
For a trace pattern $R_{i}$ in Stage-2 and threshold $H$, there must be a $j\in\{1,2,...,d\}$ such that
$f_{i}-F_{i,j,k}\geq H\Leftrightarrow F_{i,j,k}\leq f_{i}-H$
By Markov inequality, we can evaluate the probability of not recoding trace pattern $R_{i}$ with one hash function
$P(F_{i,j,k}\geq f_{i}-H)\leq\frac{E[F_{i,j,k}]}{f_{i}-H}=\frac{N-f_{i}}{m(f_{i}-H)}$
Since we have $d$ hash functions, so the retention probability of a trace pattern $R_{i}$ is
$P(R_{i})\geq 1-(\frac{N-f_{i}}{m(f_{i}-H)})^{d}$
We further conduct parameter sensitivity analysis (Section IV-D) and design space exploration (Section IV-E) to determine the optimal configuration of SL-Recorder.
III-D SL-Tracer
SL-Tracer is designed to identify and localize the root cause fail-slow failures. It follows a detection–construction–analysis pipeline. First, core-level detection and link-level detection are performed to identify potential fail-slow candidates. These candidates, together with workload’s communication dependency graph and the hardware topology, are integrated together to construct multi-level communication graph (MCG). Finally, the FailRank algorithm operates on this graph to quantify the likelihood of each candidate being the root cause, providing a ranked explanation of fail-slow components.
III-D1 Core-Level Fail-Slow Detection
The goal of core-level fail-slow detection is to identify candidate fail-slow cores. Our method performs the detection in a stage-aware manner, leveraging the mapping from DNN to many-core accelerator.
As illustrated in Figure 7, we first partition computation instructions by execution stage (e.g., network layers), ensuring that only tasks with comparable semantics are grouped together. Next, cores are clustered according to the workload-to-core mapping. Each group therefore contains volume-equivalent computations executed across different cores, making their performance directly comparable.
Within each group, we measure the performance of each core in terms of floating-point operations (FLOPs) per unit time. The group average serves as the baseline, and we perform outlier detection to highlight cores whose FLOPs are significantly lower than the baseline. These outliers are flagged as fail-slow candidates. For example, in Figure 7, core 6 and 10 deviate from their peers and are thus marked as candidates. Each candidate is further assigned a fail-slow probability, derived from its performance variance distribution. This probability is then used as the initial fail-slow score for the FailRank algorithm (Sec III-D3).
Figure 7: Core level fail-slow detection.
III-D2 Link-Level Fail-Slow Detection
The goal of link-level fail-slow detection is to identify candidate fail-slow interconnect links. Since such fail-slow behaviors cannot be directly observed under limited hardware visibility, we infer them from communication traces collected by SL-Recorder. Specifically, we formulate the problem as a statistical inference task, where observed end-to-end communication delays are decomposed to estimate the bandwidth of individual links.
For each communication event, we record four key attributes: the transferred data volume, the measured transmission time, and the index of the source and destination cores. Leveraging the deterministic routing algorithm of the 2D-mesh interconnect, we map each event to the set of links traversed along its multi-hop path, thereby associating observed end-to-end latency with individual link bandwidths.
To estimate link bandwidth, we transform the problem into a set of linear equations by turning each link’s bandwidth $b_{i}$ into its inverse $\theta_{i}=\frac{1}{b_{i}}$. Each communication event yields an equation of the form $Ax^{T}=T_{comm}$, where $A$ represents the link composition of paths, $x$ represents $\theta_{i}$ vector and $T_{comm}$ represents the observed transmission times.
Because the system is underdetermined and affected by measurement noise, we adopt the Expectation–Maximization (EM) algorithm to infer link bandwidths. EM algorithm iteratively estimates link bandwidths by alternating between (1) computing expected path-level delays given current bandwidth estimates and (2) updating bandwidth estimates to maximize the likelihood of the observed traces. This process converges to a set of link bandwidths that best explain the data, even under partial or noisy observations.
To assess the likelihood of fail-slow behavior, we calculate the fail-slow probability using the bandwidth gamma distribution model. These probabilities are then assigned as the initial fail-slow scores of links in the FailRank algorithm.
III-D3 Fail-Slow Root Cause Analysis
To localize fail-slow root causes, we construct MCG that integrates hardware topology, execution dependencies, and runtime tracing data. This augmented graph captures how performance degradation propagates across cores and links, turning the localization into a graph-based inference problem. On this foundation, we design FailRank, a PageRank-inspired algorithm that quantifies the propagation probability of fail-slows, enabling accurate root cause identification.
Figure 8: Construction of multi-level communication graph.
MCG Definition. MCG is defined as a layered graph $G=(V,E)$ that integrates task dependency and hardware topology, shown in the right of Figure 8. Each node $v\in V$ represents a core in a time window, bounded with a set of computation instructions. The likelihood for being root cause of node $v$ is quantified by a fail-slow score $s(v)\in[0,1]$ (determined in Sec III-D1). Each directed edge $e=(u,v)\in E$ denotes communication dependency between two cores, bounded with a set of communication instructions. The edge is associated with a propagation weight $w(u,v)$, capturing the probability that performance degradation propagates from core $u$ to $v$.
To construct the MCG, we perform a depth-first search (DFS) on the workload’s computation graph $G_{c}=(V_{c},E_{c})$. As illustrated in Figure 8, $G_{c}$ is a directed acyclic graph (DAG). For each dependency edge $(x,y)\in E_{c}$, we first determine the physical cores $p(x)$ and $p(y)$ assigned to tasks $x$ and $y$. The mapping $p:V_{c}\rightarrow E_{h}$ links computation nodes to hardware cores in the hardware topology $G_{h}=(V_{h},E_{h})$. The communication path between $p(x)$ and $p(y)$ is then identified as a sequence of hardware links $\{e_{1},e_{2},...,e_{k}\}\in E_{h}$. For each traversed link $e_{i}$, we increment its traffic volume $T(e_{i})$ by the transported data size of $(x,y)$.
After traversing the entire DAG, we normalize the traffic volumes to derive edge propagation weights. Specifically, for each core $u\in V_{h}$, let $Out(u)=\{(u,v)\in E_{h}\}$ denote its outgoing links. This normalization ensures that $\Sigma_{(u,v^{{}^{\prime}})\in Out(u)}w(u,v^{{}^{\prime}})=1$, making $w(u,v)$ interpretable as the probability of failure propagation through link $(u,v)$. The resulting MCG thus provides a structured, probabilistic view of how fail-slow behaviors may spread across both computation and communication dependency within a level.
To capture cross-level interactions, we introduce virtual DRAM nodes that connect levels. These nodes represent interactions through memory, such as fetches or writes between different levels. Edges connected to a DRAM node represent inter-level dependencies, showing how fail-slow may propagate across levels. This vertical connection adds a temporal dimension to MCG, allowing the FailRank algorithm to analyze the timing and causality of fail-slows.
FailRank Algorithm. To localize fail-slow root causes from core-level and link-level candidates, we propose FailRank, a graph-based iteration algorithm inspired by PageRank, which identifies fail-slow root causes on MCG. Unlike PageRank that only updates node scores based on their neighbors with uniform transition probabilities, FailRank incorporates propagation weights of their neighbor links.
The algorithm initialize fail-slow scores with the probabilities obtained in core-level and link-level detection, denoted as $s_{0}(v)$ for node $v$ and $l_{0}(u,v)$ for edge $(u,v)$. At each iteration, a node’s score is updated not only from its neighboring nodes but also from the weighted contributions of connecting links:
$s^{(k+1)}(v)=(1-\lambda)s_{0}(v)+\lambda\Sigma_{(u,v)\in E}w(u,v)s^{(k)}(u)$
where $w(u,v)$ is the normalized propagation weight of edge $(u,v)$ defined in Sec III-D3. Each link’s score is updated by three factors: the propagation weight, the fail-slow scores of its endpoint node $u$, and its current score:
$l^{(k+1)}(u,v)=\alpha w(u,v)+\beta s^{(k)}(u)+\gamma l^{(k)}(u,v)$
where $\alpha,\beta,\gamma$ are workload-dependent coefficients. Users can flexibly tune the edge update weights according to workload characteristics—for instance, setting a higher $\alpha$ for applications with intensive communication request. In our experiments, we set $\alpha=0.1,\beta=0.3,\gamma=0.6$.
Figure 9: An example of FailRank algorithm.
As Figure 9 shows, FailRank starts with initial fail-slow candidates and iteratively updates their scores until convergence. Each node is assigned a fail-slow score $s$, as annotated in the figure (e.g., $s=0.8,0.7$), which reflects the likelihood of being the root cause. Edges indicate communications between cores. Red edge has higher communication volume, while the dashed edge shows memory-access dependencies between levels. During iterations, the nodes with higher fail-slow scores and edges with heavier traffic exert stronger influence on their neighbors. After convergence, FailRank normalizes node and edge scores within each level of the MCG using a softmax function, yielding a probability distribution that pinpoints the most likely fail-slow root cause.
IV Evaluation
IV-A Experimental Setup
Evaluation Platform. We implement an event-driven simulator for many-core DNN accelerators using SimPy[52], a Python-based discrete-event simulation framework. The simulator can capture concurrent execution across thousands of processes, accurately modeling resource contention and inter-core communication. To ensure the reliability of our simulator, we validated it against the Dataflow Abstract Machine (DAM) framework[82], a parallel simulator for dataflow systems. We cross-verified total execution cycles produced by our simulator with DAM’s Communicating Sequential Processes with Time (CSPT) model on several dataflow graphs under identical workload configurations.
To faithfully capture real-world hardware behavior, our simulator incorporates statistical models[59, 29] for key performance parameters based on established distributions observed in production systems[32, 49]. For computing capacity, we model the computing capacity $P_{core}$ of each core as a normally distributed random variable: $P_{core}\sim N(\mu_{c},\sigma^{2}_{c})$ where $\mu_{c}$ represents the expected processing capacity and $\sigma^{2}_{c}$ quantifies the runtime variability. For link bandwidth, we model the transmission time $T_{link}$ using a Gamma distribution: $T_{link}\sim Gamma(\alpha,\beta)$ to capture latency variation caused by bandwidth. The shape parameter controls the asymmetry of latency distribution, while the rate parameter reflects the average link throughput.
TABLE III: Fail-slow detection accuracy.
Algorithm DarkNet-19 GoogLeNet VGG ResNet-50 Binary Tree
Acc.↑(%) FPR↓(%) Acc.↑(%) FPR↓(%) Acc.↑(%) FPR↓(%) Acc.↑(%) FPR↓(%) Acc.↑(%) FPR↓(%)
Thres 50.92(138/271) 53.29(81/152) 51.65(141/273) 54.61(83/152) 45.45(125/275) 59.09(91/154) 51.29(139/271) 56.77(88/155) 60.44(165/273) 46.05(70/152)
Mscope 68.20(193/283) 33.97(53/156) 67.14(188/280) 37.58(59/157) 68.46(191/279) 35.26(55/156) 64.47(176/273) 45.22(71/157) 79.20(217/274) 23.75(38/160)
IASO 31.65(163/515) 58.58(215/367) 37.14(156/420) 43.28(116/268) 22.53(164/728) 73.93(433/585) 13.54(221/1632) 89.85(1347/1499) 38.27(175/457) 52.20(166/318)
Perseus 66.38(235/354) 32.44(73/225) 80.75(235/291) 6.17(10/162) 45.37(235/518) 60.93(237/389) 16.84(235/1395) 87.99(1114/1266) 78.33(235/300) 11.11(19/171)
ADR 10.58(235/2222) 92.73(1941/2093) 11.17(235/2102) 92.29(1821/1973) 11.06(235/2123) 92.37(1842/1994) 11.17(235/2102) 92.29(1821/1973) 11.30(235/2079) 92.20(1798/1950)
SLOTH (Ours) 80.40(242/301) 12.64(22/174) 88.30(264/299) 13.64(24/176) 87.38(263/301) 14.61(26/178) 82.79(255/308) 16.48(30/182) 94.98(265/279) 3.18(5/157)
Workloads. We evaluate SLOTH on four representative DNN models that exhibit diverse characteristics: GoogLeNet[65], DarkNet-19[57], VGG[63], and ResNet-50[26]. These workloads span a range of architectural patterns from highly parallel branching structures to memory-intensive linear pipelines. The input dimensions of all workloads are configured to match typical throughput-oriented inference (batch size = 64). For controlled microbenchmark analysis, we additionally construct a binary tree computation graph where each node performs matrix operations. This synthetic workload isolates communication patterns from complex DNN semantics, facilitating ablation studies on probe latency and storage overhead.
The mapping configurations are generated by the Gemini framework[7], which specifies the tensor partitioning strategy and the block-to-core assignment. This deterministic mapping allows us to inject fail-slow failures at specific cores and observe their propagation through the computation graph.
Fail-Slow Dataset. We construct a fail-slow dataset comprising 152 failure instances characterized by three parameters: failure location (core-level or link-level at a 7:3 ratio), duration (uniformly distributed between 0–10 s), and slowdown rate (fixed at 10×). Failures are reused across workloads if they affect execution; those occurring on unused resources are excluded. For larger architectures (4×4, 6×6, 8×8), we generate additional failures proportional to the expanded resource count. To ensure balanced evaluation, we construct an equal number of negative samples by modifying DNN structures to produce workload traces without fail-slow failures. This dataset enables comprehensive evaluation of detection accuracy, localization precision, and robustness across diverse failure scenarios, workloads, and system scales.
Baseline. We compare SLOTH against five baseline methods originally designed for microservices, distributed storage, or server clusters. We adapt these methods to the many-core DNN accelerator domain with in-house implementations, and all the methods share the same raw trace collection infrastructure for fair comparison. For methods providing risk-level outputs, we classify moderate risk or higher as fail-slow. The parameters used by each method have been searched to match our data. Detailed description of baselines is as follows:
(1) Threshold Filtering (Thres) monitors raw performance metrics (per-core computation time, link latency) and flags components exceeding 2× the nominal latency threshold derived from workload profiling as fail-slow.
(2) Microscope (Mscope) [40] constructs a directed acyclic graph of inter-service dependencies and ranks root causes using random walk-based scoring. We model data dependencies between operators as service dependencies.
(3) IASO [55]. A peer-based detection framework that converts timeout signals into scores using AIMD combined with DBSCAN clustering. We treat each core as a peer node, using execution time violations as timeout signals.
(4) Perseus [46] builds polynomial regression models on latency-versus-throughput distributions, employing PCA and DBSCAN for outlier detection. We construct core-level model and use the 99.9th percentile as the detection threshold.
(5) Adaptive Detection at Runtime (ADR) [45] monitors both value and update frequency of performance variables using sliding windows with adaptive thresholds based on historical percentiles.
IV-B Fail-slow Root Cause Detection
We evaluate detection effectiveness across five workloads. As shown in Table III, baseline methods exhibit distinct limitations that impact their accuracy.
Thres and Mscope can detect both node and link fail-slows but suffer from low sensitivity. Thres misses 40-50% of true fail-slows across workloads, as static thresholds cannot adapt to dynamic performance variations. Mscope achieves slightly better recall but still misses 20-32% of cases, and its FPR reaches 33-54% due to incorrect attribution in complex dependency graphs. IASO, Perseus, and ADR detect more true fail-slows but are limited to node-level detection. More critically, these methods misclassify propagated slowdowns as root causes, resulting in high FPR. ADR’s adaptive thresholding suffers from the highest FPR (91-93%) by treating persistent propagated effects as independent failures, while Perseus shows workload-dependent behavior: it achieves low FPR (6.17%) on structurally simple GoogLeNet but degrade significantly on complex models like ResNet-50 (FPR: 87.99%).
Workload characteristics significantly influence detection difficulty. Binary Tree’s simple structure and regular communication patterns yield the highest accuracy across all methods. Conversely, DarkNet-19’s uniform workload mapping produces highly correlated traces, degrading inference quality—even SLOTH achieves only 80.4% accuracy here. Complex CNNs (GoogLeNet, VGG, ResNet-50) fall in between, where diverse layers provide richer trace information but also introduce more propagation paths that confuse baseline methods.
SLOTH outperforms baselines by addressing both node and link fail-slows while explicitly modeling fault propagation through iterative inference. It achieves 80-95% accuracy with low FPR (3-17%), demonstrating robust performance across different workload structures.
Figure 10: Time overhead of SL-Compiler probing.
IV-C Data Collection Time Overhead
In this experiment, we evaluate the runtime overhead of probe-based data collection under different instrumentation configurations. In the simulator, probe latency is determined by collected data volume—for example, reading the system clock incurs a 10-cycle delay. We use probe-free execution as the baseline and compare three configurations: Communication Probe (instruments communication instructions only), Computation Probe (instruments computation instructions only), and Full Probe (instruments all instructions). Each configuration is repeated multiple times to account for performance variability, with average execution time reported.
Figure 10 shows that probe-based collection introduces minimal overhead. Even under the most extensive instrumentation (Full Probe), execution time increases by at most 10% across all workloads. For targeted scenarios such as Communication Probe—suitable for network performance monitoring—overhead remains below 5%. These results demonstrate that our approach enables fine-grained trace collection with negligible performance impact.
IV-D Effectiveness of Storage Overhead Reduction
We evaluate SL-Recorder’s storage efficiency by comparing memory consumption before and after compression. The raw format records full instruction details (index, timestamps, operands, …) and serves as the baseline. For each workload, we separately measure communication and computation trace sizes, then apply compression to quantify storage savings. We also compare against IASO (communication traces), Perseus and ADR (computation traces)’ memory requirement to demonstrate SL-Recorder’s advantage.
Figure 11: Memory cost of communication traces with no optimization, IASO and SL-Recorder.
Figure 12: Memory cost of computation traces with no optimization, Perseus, ADR and SL-Recorder.
Figure 11 shows that SL-Recorder achieves significant compression compared to both raw traces and IASO. For Binary Tree, raw traces consume $10\text{\,}\mathrm{MiB}$ and IASO consume $7\text{\,}\mathrm{MiB}$, while SL-Recorder compresses to $86\text{\,}\mathrm{KiB}$—a 98.8% reduction. Similar results appear across all workloads. IASO’s algorithm only provides marginal savings (30-40% reduction) as it still retains full communication trace. Figure 12 demonstrates similar effectiveness for computation traces. Across all workloads, SL-Recorder consistently achieves 98-99% compression ratios, while Perseus and ADR achieve only 25-50% compression as they retain per-instruction records for regression analysis or adaptive threshold. Overall, SL-Recorder reduces both communication and computation trace storage by over two orders of magnitude across all workloads, significantly outperforming existing approaches and making fine-grained trace collection practical for resource-constrained environments.
Further, we examine how the internal parameters of SL-Recorder influence its compression efficiency. We conduct this experiment on DarkNet-19, varying four parameters: the number of hash functions (Hash,H), the number of buckets (Bucket,B), the Stage-2 size (Size,S), and the pattern-selection threshold (Threshold,T). To capture the interaction effects between parameters, we fix two of them at default values and vary the other two, measuring the resulting compression ratio. The results are visualized using heatmaps, where each heatmap illustrates how the compression ratio changes under different parameter combinations.
(a) S=8192,T=10
(b) B=512,T=10
(c) B=512,S=8192
(d) H=3,T=10
(e) H=3,S=8192
(f) H=3,B=512
Figure 13: SL-Recorder compression rate: Impact of the Hash (H), Bucket (B), Size (S), and Threshold (T).
Figure 13 illustrates the impact of different parameter settings on the compression ratio of SL-Recorder. Among the four parameters, H, B, and T show a clear influence on compression effectiveness, while S has little impact. Specifically, a smaller number of hash functions and buckets, as well as a larger threshold, consistently lead to stronger compression. In contrast, varying S across a wide range does not significantly alter the compression ratio, indicating that this parameter is relatively insensitive. These observations suggest that tuning hash, bucket, and threshold values is crucial for achieving high compression efficiency, whereas Stage-2 size can be fixed without noticeably affecting performance.
(a) $\times4$ Mesh.
(b) $\times8$ Mesh.
Figure 14: SL-Recorder Design Space Exploration.
IV-E Design Space Exploration
To find out the optimal configuration of SL-Recorder, we conduct a design space exploration (DSE) across four key parameters: the number of hash functions (Hash), the number of buckets (Bucket), the Stage-2 size (Size), and the pattern-selection threshold (Threshold). These parameters jointly determine the compression efficiency and detection accuracy of SLOTH.
We define a composite optimization objective as: $COST=ACC^{\alpha}\times R^{\beta}\times M^{\gamma}$, where $ACC$ represents the detection accuracy, $R$ denotes the data compression rate, and $M$ refers to the memory overhead of data structure. By adjusting the hyperparameters $\alpha,\beta,\gamma$, users can flexibly prioritize accuracy, compression rate, or memory overhead according to their own deployment requirements. Here we set $\alpha=-1,\beta=\gamma=1$, considering all three objectives equally.
Figure 14(a) shows the exploration results on a $\times4$ many-core DNN accelerator, where each point corresponds to a parameter configuration, and different colors represent different workloads. We can observe that the design space exhibits clear trade-offs: configurations with lower memory cost often suffer from degraded accuracy, whereas moderate-cost configurations yield balanced performance. The lowest-cost Pareto-optimal cases are highlighted with red triangles. Figure 14(b) shows DSE results on a $\times8$ many-core DNN accelerator, which has the same trade-off like that on $\times4$ accelerator.
Figure 15: Convergence analysis of FailRank on different fail-slow cases. The y-axis (log scale) represents the L1-norm difference between score vectors of consecutive iterations. The algorithm converges rapidly within 20 iterations for all datasets ($\epsilon=10^{-4}$).
Figure 16: Scalability of SLOTH: Time overhead.
Figure 17: Scalability of SLOTH: Memory cost and detection accuracy.
IV-F FailRank Convergence Analysis
[…]
Figure 15 illustrates the reduction of $\Delta_{k}$ over iterations on a logarithmic scale, revealing a distinct and rapid convergence trend. The process typically begins with an initial rapid descent phase during the first few iterations (iterations 1-3), where the error drops sharply by an order of magnitude as the ranking scores propagate through the immediate neighborhood of the failure. This is immediately followed by a stable decay phase, where the error decreases linearly on the log-scale plot, indicating a geometric rate of convergence as the global ranking stabilizes.
Crucially, the curves representing different fail-slow locations and intensities exhibit highly consistent trajectories, demonstrating that FailRank is robust to variations in failure patterns. In all tested cases, the algorithm satisfies the convergence criterion within approximately 17 iterations. Given that each iteration involves only lightweight matrix-vector operations, this low and bounded iteration count ensures that FailRank incurs negligible runtime overhead, making it well-suited for detection within the resource-constrained environment of many-core accelerators.
IV-G Scalability Analysis
To evaluate the scalability of SLOTH, we test its performance on many-core DNN accelerators of different scales. Specifically, we simulate ResNet-50 and DarkNet-19 on $\times4$, $\times6$, and $\times8$ accelerators, and measure the execution time under three probing modes: communication probe, computation probe, and full probe.
The results are shown in Figure 16. For both networks, the overall execution time decreases as the architecture scale grows, which indicates that larger many-core DNN accelerators could provide higher parallelism and better workload distribution. More importantly, the additional overhead introduced by probing remains consistently low across different scales. As illustrated by the red points in the figure, the normalized execution time overhead of the full probe is always below 10%, demonstrating that our framework introduces only marginal overhead even at larger scales.
[…]
V Related Work
Distributed Tracing. Existing distributed tracing systems reduce overhead through reactive collection or selective storage[34, 38, 37, 62], but these approaches are often not suitable for diagnosing fail-slow behaviors. For instance, Pivot Tracing [48] enables on-demand instrumentation based on user queries, yet critical early-stage evidence of gradual degradation is often missed. Hindsight [81] uses post-hoc triggers to commit buffered traces, but the performance decay of fail-slow issues may not activate these triggers in time, leading to the loss of crucial data. Similarly, while Sieve [31] applies an attention model to prioritize traces, it is biased towards conspicuous spikes and can overlook the weak, temporally-correlated signals characteristic of a fail-slow. In contrast, our SL-Compiler employs a proactive and structure-aware strategy; by analyzing the DNN’s computational graph to automatically instrument critical dataflows (while also allowing user guidance), it can capture most of the traces indicating fail-slow degradation.
Trace Storage Optimization. The sheer volume of trace data needs optimizations for efficient storage[43, 30, 58]. For example, Mint [30] reduces storage footprint by decomposing traces into common patterns and specific attributes. However, this structural compression is agnostic to performance semantics; it does not help in identifying or prioritizing traces that exhibit gradual latency degradation. Other works leverage sketches for data summarization[39, 75, 51]. For example, Dynamic Sketch [75] uses a doorkeeper mechanism to efficiently detect heavy hitters, but it focus on request frequency and is misaligned with fail-slow analysis. Similarly, while SketchLib [53] provides highly optimized sketch implementations for Reconfigurable Match-Action Tables (RMT), its design is incompatible with the stringent on-chip memory constraints. Our SL-Recorder employs a novel two-stage sketch design. By extending traditional sketch features to include latency statistics, it can identify similar workloads within a DNN and dynamically update them to filter the traces that are most indicative of fail-slow behavior.
Failure Detection and Root Cause Analysis. Ensuring system reliability requires robust mechanisms for both failure detection and subsequent root cause analysis (RCA)[71, 21, 20, 33]. For instance, ScalAna [33] leverages backtracking on a Program Structure Graph (PSG) to localize root causes. However, its reliance on a purely logical graph, which abstracts away the hardware topology, hinders the diagnosis of communication-related bottlenecks. Another class of methods focuses on component-level metrics[55, 46, 45]. Approaches like IASO[55], using timeout signals and peer-scoring, and ADR [45], which monitors latency and frequency shifts, are effective at identifying node-level slowdowns. Nevertheless, their node-centric perspective fails to capture performance degradation occurring on the communication links themselves. Our SL-Tracer overcomes these limitations using a Multi-level Communication Graph (MCG) that integrates software communication dependencies with the underlying physical topology. By applying an iterative FailRank algorithm on MCG, SL-Tracer can pinpoint the root cause of fail-slow, irrespective of whether the origin is a computational node or a communication link.
VI Conclusion
[…]